OpenStudy (dls):

Find the equation to the tangent to the curve x^2+3y-3=0 which is parallel to the line y=4x-5

OpenStudy (dls):

\[\LARGE \frac{dy}{dx}=2x+3=4\] \[\LARGE x=-\frac{1}{2},y=\frac{11}{12}\]

OpenStudy (dls):

\[\LARGE (y-\frac{11}{12})=4(x+\frac{1}{2})\]

OpenStudy (dls):

my attempt^

OpenStudy (amistre64):

why x=-1/2 ?

OpenStudy (dls):

oops 1/2

OpenStudy (amistre64):

other than that, your process is good

OpenStudy (dls):

12y-48x+13=0 is what im getting :O

OpenStudy (amistre64):

y = -5/4 y+5/4 = 4(x-1/2) this is a line equation ... so if form does not matter this would suffice

OpenStudy (amistre64):

otherwise, algrbrate it to your hearts content

OpenStudy (dls):

isnt y 11/12?

OpenStudy (amistre64):

.5^2 + 3(.5) - 3 .25 + 1.5 - 3 1.75 - 3 = -1.25 = -5/4

OpenStudy (amistre64):

since x = 1/2 and NOT -1/2 you have to reevaluate the y you found

OpenStudy (dls):

arent we substituting x=1/2 in the equation of curve?

OpenStudy (amistre64):

yes, which is what i just did

OpenStudy (dls):

okay! thanks! got it

OpenStudy (amistre64):

\[\frac14+\frac32-\frac31\] \[\frac14+\frac64-\frac{12}4\] \[\frac{7-12}4\]

OpenStudy (dls):

@amistre64 sorry,it was 3Y and not 3X that is how I got 11/12...

OpenStudy (dls):

it is x squared so -1/2 and 1/2 both are same that is why i didnt change the y component

OpenStudy (amistre64):

ohh then your derivative might need some looking at: x^2+3y-3=0 2x +3y' = 0 y' = -2x/3 = 4 x = 12/-2 = -6

OpenStudy (amistre64):

(6)^2+3y-3=0 33+3y=0, when y = -11

OpenStudy (amistre64):

y+11 = 4(x+6) would be the tangent line

OpenStudy (dls):

(-6,-11)

OpenStudy (dls):

CORRECT NOW! :D

OpenStudy (amistre64):

yay! ;)